Caffeine, a stimulant found in coffee and some soft drinks, consists of 49.5% C, 5.15% H, and 28.9% N by mass; the rest is oxygen. Determine the empirical formula of caffeine.
1
Expert's answer
2020-05-24T03:10:13-0400

Empirical formula

The % mass of Carbon = 49.5%

                        Hydrogen = 5.15%

                        Nitrogen = 28.9%

                    Thus, Oxygen will be (100- (49.5+5.15+28.9)) = 16.45% ÷

Assuming that the mass of the compound, caffeine, is 100g,

The mass of C is mC= 49.5g

The mass of H is mH = 5.15g

The mass of N is mN = 28.9g

The mass of O is mO= 16.45g

First, we find the number of moles of C, H, N, and O,

nC = mC ÷ MC = 49.5g ÷ 12.011g/mol = 4.121 mol

nH = mH ÷ MH = 5.15g ÷ 1.008g/mol = 5.109 mol

nN = mN  ÷ MN = 28.9g  ÷ 14.007g/mol = 2.063 mol

nO = mO ÷ MO = 16.45g ÷ 16g/mol = 1.028 mol

To get the empirical formula, we divide each number of moles with the smallest one

Smallest is 1.0281 moles

Number of C atoms = 4.121 ÷ 1.028 = 4

Number of H atoms = 5.109 ÷ 1.028 = 5

Number of N atoms = 2.063 ÷ 1.028 = 2

Number of O atoms = 1.028 ÷ 1.028 = 1

Thus, the empirical formula of caffeine is,

                                            C4H5N2O